OpenStudy (anonymous):

condense 3logd-4logf

OpenStudy (anonymous):

please help :,(

OpenStudy (agent0smith):

http://openstudy.com/study#/updates/5189ff5be4b0b0597cf8a05b see the last post

OpenStudy (hunus):

\[n*\log(x) = \log(x^n)\] And if two logs are subtracted then you divide what is in them \[\log(a)-\log(b)=\log(\frac{a}{b})\]

OpenStudy (hunus):

Do you still need help?

OpenStudy (anonymous):

no i got it thank you, sorry for the late reply, i appreciate your help guys , im going to study here more often :)

OpenStudy (hunus):

Good :D Glad we could help